An industry consists of three firms with sales of $345,000, $750,000, and $215,000.

a. Calculate the Herfindahl-Hirschman index (HHI).

Instruction: Enter your response rounded to the nearest integer.

2633 2633 Incorrect


b. Calculate the four-firm concentration ratio (C4).

26.33 + 57.25 + 16.42 + 0 + 100 26.33 + 57.25 + 16.42 + 0 + 100 Incorrect


c. Based on the FTC and DOJ Horizontal Merger Guidelines described in the text, is the Department of Justice likely to attempt to block a horizontal merger between two firms with sales of $345,000 and $215,000?

  • No

  • Yes

Please explain answers showing step by step work.
Answer #1

An industry consists of three firms with sales of $345,000, $750,000, and $215,000. Market shares are 26.34%, 57.25% and 16.41%

a. Herfindahl-Hirschman index (HHI) = squared sum of market shares

= 26.34^2 + 57.25^2 + 16.41^2

= 4241

b. Four-firm concentration ratio (C4) is the sum of market shares of top 4 firms

= 26.34 + 57.25 + 16.41

= 100%.

c. A horizontal merger between two firms with sales of $345,000 and $215,000 will result in two firms in the market with shares 57.25% and 42.75%. New HHI is 57.25^2 + 42.75^2 = 5105. Since the HHI is increased by more than 100, it will be blocked

Yes.